# Direction Estimation of Sound Sources Using Microphone Arrays and Signal Strength

### v2

@@ Abstract
Index Terms:
Sound tracking, Sound source localization, Direction estimation, Microphone arrays
The source code of the project is available on our GitHub repository:
https://github.com/mahdi943/soundlocalization
I Introduction
@@
While sound localization generally refers to estimating the exact position of a sound source in space, sound tracking specifically focuses on determining the direction from which the sound originates. Localization aims to provide precise spatial coordinates, which often requires complex sensor arrays and extensive computation. In contrast, sound tracking emphasizes directional information, enabling real-time responsiveness with fewer sensors and simpler processing. This distinction allows sound tracking to serve as a lightweight, cost-effective alternative in scenarios where directionality is sufficient for the task at hand.
Many existing sound localization methods rely on complex hardware setups, intensive signal processing, and high computational demands. While these approaches offer high accuracy, they are often impractical for lightweight, low-cost applications. To address this, we propose a simple and efficient method for detecting the direction of sound sources using a minimal sensor setup, designed to operate with limited processing resources.
This work focuses on enabling directional sound-tracking for use in responsive systems such as smart home devices, interactive robots, and compact automation solutions. By estimating the direction of incoming sounds in real time, the proposed method supports adaptive behavior in resource-constrained environments, offering a practical alternative to more computationally intensive localization techniques.
Our approach employs a vector-based technique to accurately detect sound direction by computing the sum of sound signal vectors arranged in a rectangular configuration. For instance, if a sound originates from the left, the vector representation will reflect this directional characteristic. This capability is crucial for swiftly identifying potential threats in security systems and enhancing the responsiveness of smart home devices by dynamically adjusting settings based on user positions.
While machine learning and beamforming techniques are prevalent in the field of sound localization, our method distinguishes itself through its simplicity and affordability. By utilizing readily available Arduino boards and electret microphones, we simplify the assembly and configuration process, thereby enhancing accessibility for various applications. This approach holds great promise for improving systems that rely on precise sound source detection and tracking, especially in environments where budget constraints or deployment speed are critical considerations.
II Related Work
II-A Sound Source Localization (SSL)
@@
II-B Sound Source Tracking
@@ II-C Summary and Motivation
In contrast, our work focuses on a simple, intensity-based method that relies on only three microphones and does not require time delay or phase computations. This enables a more practical, scalable, and cost-effective solution for indoor directional sound-tracking, suitable for applications such as smart homes, speaker-following systems, and acoustic event detection.
III Approach
In this section, we describe the hardware and software setup used to implement our sound localization system. We then provide a detailed explanation of the components, the underlying methodology, and the measurement process involved in evaluating the system’s performance.
III-A Methodology
@@
1 Input: Sound signal $S$, number of samples NN, threshold $T$
@@ 24 $Servo\leftarrow ServoPos$
These polar vectors are converted to rectangular (Cartesian) form using standard trigonometric conversion, resulting in vectors $R=(x,y)$, where $x=r\cdot\cos(\theta)$ and $y=r\cdot\sin(\theta)$.
@@
The three rectangular vectors are summed component-wise $R_{\text{total}}=(X,Y)$ to compute the resulting direction vector, where:
@@
The final direction of the sound source is determined by converting the resultant rectangular vector back to polar coordinates. The magnitude and angle are computed as follows:
@@
III-B Implementation
The implemented sound-tracking system consists of three electret microphones arranged on a circular surface at equal angular intervals (0°, 120°, and 240°), each positioned 15 cm from the center. These microphones capture sound signals from different directions and are connected to LM358 operational amplifiers to boost the signal strength. The amplified outputs are fed into an Arduino board, which processes the signals and determines the direction of the incoming sound. A servo motor, also connected to the Arduino, rotates an indicator shaft toward the detected sound direction. Figure 2 illustrates the complete hardware setup, which is designed to be low-cost, efficient, and suitable for compact applications.
@@ III-B Implementation
IV Results
IV-A Experimental Setup and Design
@@
IV-B Vector Visualization and Interpretation
@@
Figure 4: Detecting sound source angles and magnitude at 0 degrees
@@
IV-C Accuracy and Precision Analysis
@@
Table I: Accuracy and Precision at Different Angles
@@
IV-D Conclusion
Overall, the results demonstrate that the proposed sound-tracking system is capable of estimating the direction of a sound source with high accuracy and consistency. The system remains robust under different angular positions and moderate environmental noise, making it suitable for applications such as smart home automation and security surveillance.
V Future Work
Based on the method presented in this study, there are several potential routes for future work to improve and expand the sound-tracking system’s capabilities. One possible enhancement would be to improve the precision of the sound-tracking by including more accurate microphones and implementing a beam-forming technique in addition to our approach to compute the direction of the sound source. Furthermore, incorporating machine learning techniques could improve the system’s ability to classify the direction of the sound based on audio signal characteristics, as well as anticipate the distance of the sound source.
Another direction for future work is to expand the capabilities of the system to track multiple sounds simultaneously by using multiple servo motors, each controlled by a separate set of microphones and vectors. Developing a user interface to display the direction of the sound and control the servo motor in real time could also be a valuable addition to the system.
To increase robustness, future versions of the system could integrate noise reduction methods and microphone calibration to mitigate environmental interference. The low-cost and low-complexity design has proven effective in controlled tests, and extending this to more complex environments—such as those with physical or acoustic obstacles—could help assess the system’s real-world performance and limitations.
Testing the system in various environments and scenarios may provide useful insights into its performance and highlight opportunities for improvement. Exploring the method’s possible applications in noise-canceling devices or security systems could also open up new avenues for this technology.
VI Conclusion
This paper presents a simple and cost-effective method for sound-tracking using an Arduino, three electret microphones, three LM358 op-amps, and a servo motor. By calculating the sum of vectors in rectangular form, we were able to determine the direction of the sound source with high accuracy. The results of our experiment demonstrated that the servo motor was able to track the direction of the sound source, and the recorded audio was clear and reliable.
The technique we presented in this study offers a practical and affordable solution for sound localization and tracking, with potential applications in various areas such as security, surveillance, and audio recording. It is particularly useful in scenarios where directional audio is required, such as in video conferencing, sound engineering, and audio equipment testing.
Direction detection plays a critical role in sound source localization, as it allows systems to spatially orient themselves toward a sound-emitting object. This directional awareness enhances the system’s ability to respond to audio stimuli, improves context awareness in robotics and automation, and enables more efficient resource focus in smart environments. Accurately detecting the direction from which a sound originates is often the first step in identifying, isolating, or responding to the source—making it a cornerstone for real-world applications.
Moreover, this method provides an excellent alternative to expensive commercial sound-tracking systems, which are often complex and difficult to use. It offers a simple and effective way to detect and record audio and control a servo motor based on the direction of the sound.
Finally, our study provides a valuable contribution to the field of sound-tracking and offers a promising solution for sound localization and tracking with an affordable and practical approach. Future research could focus on improving the design and implementation of the sound-tracking system by incorporating additional microphones and sensors to enhance its accuracy and precision. Furthermore, additional testing and experimentation could be conducted to evaluate the system’s performance in different environments and conditions.

### v3

@@ Abstract
The source code of the project is available on our GitHub repository: https://github.com/mahdi943/soundlocalization
1 Introduction
@@
While sound localization estimates the exact position of a sound source, sound tracking focuses on determining its direction. Localization typically requires complex sensor arrays and heavy computation, whereas tracking provides directional information in real time with fewer sensors and simpler processing, offering a lightweight, cost-effective alternative when full coordinates are unnecessary.
Many existing methods rely on complex hardware and intensive signal processing, which limits their practicality for low-cost applications. Our approach demonstrates that reliable direction estimation can be achieved using inexpensive, widely available hardware, without high-end processors or elaborate sensor arrays. This makes it suitable for educational, prototyping, and resource-constrained environments.
This work focuses on enabling directional sound-tracking for responsive systems such as smart homes, robots, and compact automation tools. By estimating sound direction in real time, the method supports adaptive behavior in resource-constrained environments as a lightweight alternative to computationally intensive localization techniques.
Our approach uses a vector-based method to compute the sum of microphone signal vectors, allowing accurate direction detection with minimal hardware. This capability is valuable for security, automation, and smart devices that benefit from rapid and reliable sound tracking.
Unlike machine learning or beamforming approaches, our method emphasizes simplicity and affordability. Using readily available Arduino boards and electret microphones, we lower hardware costs and configuration effort, making sound-tracking accessible for budget-sensitive or rapid-deployment applications.
2 Literture Review
2.1 Sound Source Localization (SSL)
@@
In contrast to these hardware-heavy or computationally demanding approaches, our work emphasizes a lightweight solution that leverages low-cost, readily available components. This novelty allows us to achieve directional estimation without the complexity and resource requirements of traditional methods.
2.2 Sound Source Tracking
@@ 2.3 Summary and Motivation
Instead, our work focuses on a simple, intensity-based method that relies on only three microphones and does not require time delay or phase computations. This enables a more practical, scalable, and cost-effective solution for indoor directional sound-tracking, suitable for applications such as smart homes, speaker-following systems, and acoustic event detection.
3 Methodology
@@
1Input: Sound signal $S$, number of samples NN, threshold $T$
@@ 24 $Servo\leftarrow ServoPos$
These polar vectors are converted by Equations 4, 5 and 6 to rectangular (Cartesian) form using standard trigonometric conversion, resulting in vectors $R=(x,y)$, where $x=r\cdot\cos(\theta)$ and $y=r\cdot\sin(\theta)$.
@@
The three rectangular vectors are summed component-wise $R_{\text{total}}=(X,Y)$ to compute the resulting direction vector, as defined in Equations 7 and 8.
@@
The final direction of the sound source is determined by converting the resultant rectangular vector back to polar coordinates. The magnitude and angle are computed as shown in Equations 9 and 10.
@@
3.1 Implementation
The implemented sound-tracking system consists of three electret microphones arranged on a circular surface at equal angular intervals (0°, 120°, and 240°), each positioned 15 cm from the center. These microphones capture sound signals from different directions and are connected to LM358 operational amplifiers to boost the signal strength. The amplified outputs are fed into an Arduino Uno R3 microcontroller, which processes the signals and determines the direction of the incoming sound. A servo motor, also connected to the Arduino, rotates an indicator shaft toward the detected sound direction. Figure 2 illustrates the complete hardware setup, which is designed to be low-cost, efficient, and suitable for compact applications. The low sampling rate (50 Hz) necessitated placing the sound source close to the microphone array (35 cm) to ensure reliable direction estimation. While this limits the testing range, it demonstrates the feasibility of ultra-low-cost hardware.
4 Results
4.1 Experimental Setup and Design
@@
4.2 Vector Visualization and Interpretation
@@
Figure 4: Detecting sound source angles and magnitude at 20 degrees
@@
4.3 Accuracy and Precision Analysis
@@
Table 1: Accuracy and Precision at Different Angles
@@
5 Conclusion
This paper introduced a simple and cost-effective method for sound-tracking using an Arduino, electret microphones, LM358 op-amps, and a servo motor. By applying a vector-sum approach, the system successfully estimated sound direction and controlled the servo to track the source. The results highlight the potential of this lightweight setup for practical applications in security, surveillance, robotics, and smart environments where directional audio is essential.
The proposed method demonstrates an affordable alternative to complex and expensive sound-tracking systems, making it accessible for real-time use in constrained environments. While effective, the experiments were limited by the Arduino Uno’s low sampling rate and memory capacity, requiring the sound source to remain close to the microphone array. This design choice reflects our focus on validating the algorithm’s feasibility on low-resource hardware rather than optimizing for large-scale deployment.
Future work will address these limitations by exploring higher sampling rates, more capable microcontrollers, and additional microphones or sensors to improve accuracy. Testing under realistic, noisy conditions will further evaluate robustness, extending the method’s applicability. Overall, this study contributes a practical foundation for affordable sound-tracking and opens opportunities for enhanced systems in robotics, automation, and acoustic sensing.